A massive steel cable drags a 20 kg block across a horizontal, frictionless surface. A 90 N force applied to the cable causes the block to reach a speed of 4.0 m/sin a distance of 2.2 m .
Answer #1

here

a = v^2 / 2 * x

a = 4 * 4 / 2 *2.2

a = 3.64 m/s^2

then

F = m * a

90 = (20 + m1) * 3.64

m1 = 4.52 kg